# MedReasoner: Reinforcement Learning Drives Reasoning Grounding from Clinical Thought to Pixel-Level Precision

### v1

@@ A.2 Language-Guided Medical Image Segmentation
To support interactive and context-sensitive segmentation, some methods generate language prompts dynamically and enforce geometric consistency. FLanS (Da et al. 2024) uses retrieval-augmented prompts with a geometry-aware model for orientation-consistent segmentations. Besides, CLIP-based models (Liu et al. 2023) inject general language priors to support zero-shot anatomical segmentation.
@@
Another line of research enhances LGMIS by combining pretrained vision-language and segmentation models under weak or zero-shot supervision. MedCLIP-SAM (Koleilat et al. 2024) fuses BiomedCLIP and SAM with ScoreCAM-based supervision to achieve weakly supervised yet accurate segmentation.
@@ A.3 Referring Lesion Segmentation
These approaches aim to enhance the alignment between clinical language and visual features, often through contrastive learning, causal reasoning, or language-conditioned modeling. Li et al. (Li et al. 2024a) apply multi-level contrastive learning, while CausalCLIPSeg (Chen et al. 2024b) introduces causal interventions to suppress spurious correlations. LViT (Li et al. 2023b) incorporates language-guided attention into transformers for holistic lesion understanding.
@@
This direction focuses on extending the capabilities of pretrained segment models to the medical domain with minimal supervision. LGA (Hu et al. 2024) enables controllable segmentation by injecting language cues into SAM via a lightweight adapter. SimTxtSeg (Xie et al. 2024) and (Huang et al. 2024) utilize prompt-based pseudo-labeling and language-conditioned reconstruction for weakly supervised alignment.
@@ B.1 Meta Information of U-MRG-14K
Following the established methodologies for constructing diverse datasets (Qiao et al. 2024; Song et al. 2025; Diao et al. 2024, 2025; Wang et al. 2025a, b), we design and curate U-MRG-14K with a focus on medical image reasoning across multiple modalities and hierarchical categories.
@@ Appendix C More Details on MedReasoner Framework
We utilize Group Relative Policy Optimization (GRPO) (Shao et al. 2024) as our reinforcement learning strategy. is an efficient reinforcement learning algorithm that eliminates the need for a value network by leveraging group-wise relative advantages. For each input query $q$, GRPO samples a group of $G$ outputs $\{o_{i}\}_{i=1}^{G}$ from the old policy $\pi_{\text{old}}$, scores them using a reward model $r_{\phi}$, and computes normalized relative advantages within the group. The policy $\pi_{\theta}$ is updated by maximizing the following objective:
@@ D.1 Implementation Details
We adopt Lingshu-7B (Xu et al. 2025) and MedSAM2 (Ma et al. 2025) as our default CRM and ASM, respectively. MedReasoner is trained on an 8 NVIDIA A100-80G GPUs with the veRL (Sheng et al. 2025). Training utilized a total batch size of 5 and 16 samples per step. The initial learning rate is set to 1e-6. In terms of generation configuration, we follow the same settings as used in Lingshu, setting the repetition penalty to 1.05 and the temperature to 0.1, while employing argmax sampling.
@@ D.2 Model Details
Gemini 2.5 Flash (Google 2025)

### v2

@@ A.2 Language-Guided Medical Image Segmentation
To support interactive and context-sensitive segmentation, some methods generate language prompts dynamically and enforce geometric consistency. FLanS (da2024segment) uses retrieval-augmented prompts with a geometry-aware model for orientation-consistent segmentations. Besides, CLIP-based models (liu2023clip) inject general language priors to support zero-shot anatomical segmentation.
@@
Another line of research enhances LGMIS by combining pretrained vision-language and segmentation models under weak or zero-shot supervision. MedCLIP-SAM (koleilat2024medclip) fuses BiomedCLIP and SAM with ScoreCAM-based supervision to achieve weakly supervised yet accurate segmentation.
@@ A.3 Referring Lesion Segmentation
These approaches aim to enhance the alignment between clinical language and visual features, often through contrastive learning, causal reasoning, or language-conditioned modeling. Li et al. (li2024language) apply multi-level contrastive learning, while CausalCLIPSeg (chen2024causalclipseg) introduces causal interventions to suppress spurious correlations. LViT (li2023lvit) incorporates language-guided attention into transformers for holistic lesion understanding.
@@
This direction focuses on extending the capabilities of pretrained segment models to the medical domain with minimal supervision. LGA (hu2024lga) enables controllable segmentation by injecting language cues into SAM via a lightweight adapter. SimTxtSeg (xie2024simtxtseg) and (huang2024cross) utilize prompt-based pseudo-labeling and language-conditioned reconstruction for weakly supervised alignment.
@@ B.1 Meta Information of U-MRG-14K
Following the established methodologies for constructing diverse datasets (qiao2025we; song2024cs; diao2025seas; diao2025driverx; wang2025harnessing; wang2025cinetechbench), we design and curate U-MRG-14K with a focus on medical image reasoning across multiple modalities and hierarchical categories.
@@ Appendix C More Details on MedReasoner Framework
We utilize Group Relative Policy Optimization (GRPO) (shao2024deepseekmath) as our reinforcement learning strategy. is an efficient reinforcement learning algorithm that eliminates the need for a value network by leveraging group-wise relative advantages. For each input query $q$, GRPO samples a group of $G$ outputs $\{o_{i}\}_{i=1}^{G}$ from the old policy $\pi_{\text{old}}$, scores them using a reward model $r_{\phi}$, and computes normalized relative advantages within the group. The policy $\pi_{\theta}$ is updated by maximizing the following objective:
@@ D.1 Implementation Details
We adopt Lingshu-7B (xu2025lingshu) and MedSAM2 (ma2025medsam2) as our default CRM and ASM, respectively. MedReasoner is trained on an 8 NVIDIA A100-80G GPUs with the veRL (sheng2025hybridflow). Training utilized a total batch size of 5 and 16 samples per step. The initial learning rate is set to 1e-6. In terms of generation configuration, we follow the same settings as used in Lingshu, setting the repetition penalty to 1.05 and the temperature to 0.1, while employing argmax sampling.
@@ D.2 Model Details
Gemini 2.5 Flash (gemini2.5_flash)